Find the general value of ii…........̣.........................................

Grade:11

1 Answers

Aditya Gupta
5 years ago
the principle value of i^i is exp(-Pi/2) which is approximately 0.20788.
this is bcoz
i=e^ipi/2
so i^i= e^i2pi/2=e^ –pi/2